OpenStudy (anonymous):

PLEASE HELP ME IDK HOW TO WRITE THESE PLEASE HELP!!!!!!! (During a clothing store's Bargain pays, the regular price for T-shirts is discounted by $5 there us a state tax of %5 and the $5 discount is applied before the sales tax is calculated. a.write an expression that shows the regular price r of a T-shirt minus the $5 discount. b.Write a rule for the function p(r) that expresses the final price p of a T-shirt with discount applied and sales tax added. c.how much would you pay during Bargain Days for a shirt regularly priced at $15.50.

OpenStudy (math&ing001):

a. price = regular price-$5 b. P(r)=(r-5)+(r-5)*5/100=(r-5)1.05 c. calculate P(r) with r=$15.50

OpenStudy (anonymous):

Thank you omg ur a lifesaver(:

OpenStudy (math&ing001):

Anytime

OpenStudy (anonymous):

i have another problem do u think u can help me with it? its my last one i promise